Two projectiles are thrown with the same initial velocity, one at an angle and the other at


an angle of . (a) Can both projectiles strike the ground at the same distance from the


projection point? (b) Can both projectiles be in air for the same time interval?


1
Expert's answer
2022-02-23T08:31:55-0500

a)

YES

"\sin2\alpha=\sin2(90°-\alpha),"

b)

NO

"\sin\alpha \not =\sin(90°-\alpha),"

YES, when "\alpha=45°."